OpenStudy (autumn9913):

From a phone survey, a researcher determined that the true estimate for the proportion of voters who will reelect the mayor is likely to be the interval (0.39, 0.47). What is the survey's margin of error?

OpenStudy (michaelbp):

The margin of error from a survey is half the witdth of the interval. Margin of error = width of the interval / 2 = (0.47 - 0.39) / 2 = 0.08 / 2 = 0.04